Three equal charges of +10 nC are located at the c
Phys 2: steps are much appreciated, thanks!
Answer #1

10nC Q2 Q1 10nC

Calculate the forces between any two charges

\large F=\frac{1}{4\pi \epsilon _0}\frac{Q_1Q_2}{r^2}

\large F=k\frac{Q^2}{r^2}   as we know here \large Q_1=Q_2=Q=10nC=10\times 10^{-9}C

\large k=9\times 10^9Nm^2C^{-2}

Two forces are applied at each corner with a \large 60^0 angle between them

the resultant force at each corner,

\large F_{Resultant}= 2Fcos30 ^0

\large F=\frac{9\times 10^9\times (10\times 10^{-9})^2}{(10\times 10^{-2})^2}

\large F=\frac{9\times 10^9\times (100\times 10^{-18})}{0.01}=9.0 \times 10^{-5}N

\large F_{Resultant}= 2\times 9.0\times 10^{-5}cos30 ^0

\large F_{Resultant}= 15.588\times 10^{-5}N